OpenStudy (anonymous):

Suppose F(x) = the integral of: sin (t^2+1) as x goes from x^2 to e^(2x) dt Evaluate F'(x).

OpenStudy (anonymous):

\[F(x)=\int_{x^2}^{e^{2x}}sin(t^2+1) dt\]?

OpenStudy (anonymous):

i think gimmick is to break it up into two pieces. for example if you want the derivative of \[F(x)=\int_0^{e^{2x}}\sin(t^2+1)dt\] by the fundamental theorem and the chain rule you get \[F'(x)=\sin(e^{4x}+1)2e^{2x}\]

OpenStudy (anonymous):

and \[F(x)=\int_{x^2}^0 sin(t^2+1)dt=-\int_0^{x^2}sin(t^2+1)dt\]\]

OpenStudy (anonymous):

whose derivative is \[F'(x)=-\sin(x^2+1)2x\]

OpenStudy (anonymous):

oh look! i see trees!

myininaya (myininaya):

my thingy got messed up

OpenStudy (anonymous):

\[\color{green}{\text{trees are green}}\]

OpenStudy (anonymous):

\[\color{green}{\text{sorry about your thingy}}\]

myininaya (myininaya):

\[F(x)=G(e^{2x})-G(x^2) \] where G represents the antiderivative of \[g(t)=\sin(t^2+1)\] \[F'(x)=(e^{2x})'g(e^{2x})-(x^2)'g(x^2)\]

OpenStudy (anonymous):

answer is \[F'(x)=2e^{2x}\sin(e^{4x}+1)-2x\sin(x^4+1)\]

myininaya (myininaya):

right!

myininaya (myininaya):

i c no trees here

OpenStudy (anonymous):

\[\color{#FF0099}{\text{time for you to get an icon of some sort}}\]

OpenStudy (anonymous):

\[\color{#00FF11}{\text{no self respecting superhero should be a chair with feet}}\]

OpenStudy (anonymous):

\[\color{red}{\text{btw sorted mathmom for you while you went to eat}}\]